OpenStudy (anonymous):

Out of 5 cards, 3 are blue and 3 are red. What is the probability of drawing a blue card, replacing it, and then drawing a blue card

terenzreignz (terenzreignz):

Please check your question... 3 blue and 3 red make six cards! :P

terenzreignz (terenzreignz):

Unless there's a card that's both blue and red though... hmm

OpenStudy (anonymous):

2 reds/ -.- and 3 blues.

terenzreignz (terenzreignz):

OH ok :D First, what's the probability that you draw a blue card (just once) :)

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

right?

terenzreignz (terenzreignz):

No :) Think of it this way There are 5 cards, 3 of which are blue.... and 2 of which are red It should mean that you're more likely to pick a blue card, right?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

But by how much? For instance, what are the odds of drawing a red card?

OpenStudy (anonymous):

2/5

terenzreignz (terenzreignz):

But you said the probability of getting a blue card was 1/5... something must have been wrong there ;)

terenzreignz (terenzreignz):

so what are the odds of getting a blue card? :)

OpenStudy (anonymous):

i don't know, that's why i asked you. :X lol

terenzreignz (terenzreignz):

Well, you did say that the odds of getting red are 2/5... how did you get that? :)

OpenStudy (anonymous):

so the answer is 3/5ths, right?

terenzreignz (terenzreignz):

That's right :) Now, whatever you pick the first time, it won't affect your pick for the second draw, right?

OpenStudy (anonymous):

yepp

terenzreignz (terenzreignz):

They're what you call "independent" events, and the odds of independent events happening are simply the product of their individual probabilities... did you get that? :)

OpenStudy (anonymous):

yes i did.

terenzreignz (terenzreignz):

So, the first event's probability is 0.6 and the second event's probability is also 0.6, since you return the first card Therefore, the probability that they both will happen is...?

terenzreignz (terenzreignz):

Well, if you got \[(\frac{3}{5})(\frac{3}{5})= \frac{9}{25} = 36\%, \]You're right!